OpenStudy (anonymous):

multiply z2/z+1.z2+3z+2/z2+3z

OpenStudy (anteater):

\[(\frac{ z^2 }{z+1 ? })(\frac{ z^2+3z+2 }{z^2+3z ? })\] I want to make sure I am reading it correctly. Does it look like that (what I just typed)?

OpenStudy (anteater):

Well, except for the question marks - don't know how those got in there!

OpenStudy (anteater):

If so you could factor some of the expressions and then cancel the common factors out to simplify.

OpenStudy (anteater):

For example, z^2 + 3z = z(z + 3) and z^2 = z(z), so you could cancel out a z

OpenStudy (anteater):

So you can factor the expressions and do the multiplication, then check against this if that helps. :) 

OpenStudy (anonymous):

no ( ) or ? marks but yes